OpenStudy (anonymous):

Simplify the expression below. (4x)3

hartnn (hartnn):

use this : \(\huge (ab)^n=a^nb^n \)

hartnn (hartnn):

and you know 4^3= ?

OpenStudy (anonymous):

okay so do i times 4 and 3 ? @hartnn im not good in math really would like your help

hartnn (hartnn):

nopes, in a^b you multiply a 'b' times, so in 4^3 you multiply 4 3 times, what u get ?

OpenStudy (anonymous):

64

hartnn (hartnn):

yup, 64 is correct :) so what will be (4x)^3 ?

OpenStudy (anonymous):

okay so it will be 64x^3

hartnn (hartnn):

that is correct! good :)

OpenStudy (anonymous):

okay here's another let me do it first you tell me if its wrong . (3x)2

OpenStudy (anonymous):

9x^2

hartnn (hartnn):

that is correct!

OpenStudy (anonymous):

this one different 9x(5x2)

OpenStudy (anonymous):

it should be ^2

hartnn (hartnn):

that 'x' in middle is variable \(x\) or multiplication?

OpenStudy (anonymous):

variable

OpenStudy (anonymous):

9x(5x^2)

hartnn (hartnn):

ok, so group like terms, like 9*5 and x*x^2 whats x*x^2 = ?

OpenStudy (anonymous):

its 45^2

hartnn (hartnn):

9*5 = 45, this is correct. but what is x*x^2= ? \(\huge x^n.x^m=x^{m+n}\)

OpenStudy (anonymous):

9*2 is what your asking ?

hartnn (hartnn):

nopes, \(\large 9x(5x^2) = (9*5) (x*x^2)=45x^{1+2}=?\)

OpenStudy (anonymous):

oh so 45^3

hartnn (hartnn):

45x^3 did u get how ?

OpenStudy (anonymous):

yes x+x will get you two . but this question here